OpenStudy (anonymous):

Kevin wants to make a mixture that is 50% lemon juice and 50% lime juice. How much 100% lemon juice should he add to a juice that is 40% lemon juice and 60% lime juice to make 6 gallons of the 50% lemon/50% lime juice mixture? (Points : 1)

OpenStudy (anonymous):

V = 6 quarts just focus on the lemon content 0.5 V = 3 = x + 0.4 (6-x) x = quarts of 100% lemon juice 3 = x + 2.4 - 0.4 x 3 = 0.6 x + 2.4 0.6 = .6 x x = 1 check 1 + 0.4(5) = 3 qts lemon in 6 qts total thus 50% lemon

OpenStudy (anonymous):

Help m!!!!!!!

OpenStudy (anonymous):

Nancy bought 360 tropical fish for a museum display. She bought 8 times as many goldfish as angelfish. How many of each type of fish did she buy? Which is a system of equations to model the problem if x represents the number of angelfish Nancy bought, and y represents the number of goldfish she bought? (Points : 1)